OpenStudy (anonymous):

What are the foci of the ellipse given by the equation (x-5)^2/81 +(y+1)^2/9=1 ??

OpenStudy (anonymous):

deperate need of help

OpenStudy (anonymous):

find d eccentricity 1st

OpenStudy (anonymous):

standard form of an ellipse....x^2/a^2 + y^2/b^2=1

OpenStudy (anonymous):

b^2= a^2( 1- e^2)

OpenStudy (anonymous):

e= eccentricity

OpenStudy (anonymous):

does it help u?

OpenStudy (anonymous):

bt the centre of d ellipse is not (0,0)...so there'll be some change in finding the focii

OpenStudy (anonymous):

can u find d eccentricity using above formula?